A 2100 kg pickup truck collided with a stopped car, which had a mass of 1200 kg. The two vehicles sink into each other and travel together at a speed of 4.50 m / s. The initial speed of the van is ____ m / s.


1
Expert's answer
2021-01-12T12:15:06-0500

From the law of conservation of momentum, we have:


"m_1v_{1i}=(m_1+m_2)v_f,""v_{1i}=\dfrac{(m_1+m_2)v_f}{m_1},""v_{1i}=\dfrac{(2100\ kg+1200\ kg)\cdot 4.50\ \dfrac{m}{s}}{2100\ kg}=7.1\ \dfrac{m}{s}."

Answer:

"v_{1i}=7.1\ \dfrac{m}{s}."